OpenStudy (anonymous):

show that integral x / (x^2 + 2 ) ^2 converges on [ 0,oo), use integral comparison

OpenStudy (anonymous):

probably can compare it to \[\int_0^{\infty}\frac{x}{x^4}dx\] since it is smaller

OpenStudy (anonymous):

no doesnt work, that diverges

OpenStudy (anonymous):

at least for the infinity part

OpenStudy (anonymous):

integral 1/ x^3 diverges on [ 0, oo)

OpenStudy (anonymous):

this is not an improper integral on the left, because it is finite at 0.

OpenStudy (anonymous):

in fact it is 0 at 0

OpenStudy (anonymous):

you mean the integrand is zero at zero , integral f(x) , f(0) = 0

OpenStudy (anonymous):

right. the integrand

OpenStudy (anonymous):

x/ (x^2 + 2 ) ^2 is zero when x = 0 but the comparison wont work

OpenStudy (anonymous):

sorry that is not what i mean

OpenStudy (anonymous):

i mean to say \[\int_0^1\frac{x}{(x^2+2)^2}dx\] is finite

OpenStudy (anonymous):

and \[\int_1^{\infty}\frac{x}{(x^2+2)^2}dx\] converges by comparison test

OpenStudy (anonymous):

that other stuff i wrote was nonsense, sorry

OpenStudy (anonymous):

in fact i think you can just go ahead and compute this integral if you like, because you can use a simple u - sub to get the anti derivative

OpenStudy (anonymous):

it is \[-\frac{1}{2(x^2+2)}\] at 0 you get \[\frac{-1}{4}\] and as r goes to infinity you get 0

OpenStudy (anonymous):

so integral is in fact just \[\frac{1}{4}\]

OpenStudy (anonymous):

so if it wasnt easily integrable, you would have to split the integral